In a national basketball​ association, the top​ free-throw shooters usually have probability of about 0.90 of making any given free throw. Complete parts a through c.

a. During a​ game, one such player shot 10 free throws. Let X=number of free throws made. What must you assume in order for X to have a binomial​ distribution?

A. It is assumed that the data are not binary.

B.It is assumed that the data are​ binary, that probabilities of success for trials​ (free throws) differ for each​ trial, and that the trials are independent.

C. It is assumed that the data are​ binary, that there is the same probability of success for each trial​ (free throw), and that the trials are dependent.

D.It is assumed that the data are​ binary, that there is the same probability of success for each trial​ (free throw), and that the trials are independent.

b. Specify the values of n and p for the binomial distribution of X in part a.

n

=

1010

p

=

. 90.90

​(Type integers or decimals. Do not​ round.)

c.Find the probability that the player made all 10 free​ throws, 9 free​ throws, and more than 7 free throws. The probability that the player made all 10 free throws is 0.3870.387.

​(Round to the nearest thousandth as​ needed.)

The probability that he made 9 free throws is 0.3870.387.

​(Round to the nearest thousandth as​ needed.)

The probability that he made more than 7 free throws is 0.9470.947. ​(Round to the nearest thousandth as​ needed.)
Answer #1

solution - (a) het X = number made Then of free throws the assumptions distribution for to a binomial are D) It is assumed thmore given Prob that he made them 7 free throws is by Plx>77 = P(x=8] +P/x=8] + P(x=10] 10Gq (0.9)% (0.1) + 0.349 + 0-387 = 4any query, please ask